xx Constants

Constants

Frank Lxffler (knarf@cct.lsu.edu)

Date

Abstract

Thorn to provide macros for commonly used constants like the speed of light, the gravitational constant or the solar mass.

1 Comments Because the number of defines could grow, I define some naming conventions:

Names have to start with CONSTANTx.

Afterwards a typical name is following. A x is coming next. Either SI, cgi or C (for Cactus) is the last part. This specifies the system of units used here (SI, cgi or Cactus units).

2 Provided Constants

: the speed of light : the gravitational constant : the solar mass

3 Parameters

4 Interfaces General Implements: constants

Adds header: constants.h

5 Schedule This section lists all the variables which are assigned storage by thorn EinsteinBase/Constants. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE